OpenStudy (cherie_magee):

Which of the following equations could be the result of using the comparison method to solve the system shown? x + 2y = 6 x - 4y = 8 2y - 6 = 4y + 8 6 - 2y = -4y + 8 6 - 2y = 4y + 8

OpenStudy (anonymous):

x+2y=6 =>x=6-2y so now if we replace the value of x in another equation we get 6-2y-4y=8 =>6-2y=4y+8